### v1

# DFPE: A Diverse Fingerprint Ensemble for Enhancing LLM Performance

Abstract
Large Language Models (LLMs) have shown remarkable capabilities across various natural language processing tasks but often struggle to excel uniformly in diverse or complex domains. We propose a novel ensemble method - Diverse Fingerprint Ensemble (DFPE), which leverages the complementary strengths of multiple LLMs to achieve more robust performance. Our approach involves: (1) clustering models based on response "fingerprints" patterns, (2) applying a quantile-based filtering mechanism to remove underperforming models at a per-subject level, and (3) assigning adaptive weights to remaining models based on their subject-wise validation accuracy. In experiments on the Massive Multitask Language Understanding (MMLU) benchmark, DFPE outperforms the best single model by 3% overall accuracy and 5% in discipline-level accuracy. This method increases the robustness and generalization of LLMs and underscores how model selection, diversity preservation, and performance-driven weighting can effectively address challenging, multi-faceted language understanding tasks.
DFPE: A Diverse Fingerprint Ensemble for Enhancing LLM Performance
Seffi Cohen                        Niv Goldshlager                        Nurit Cohen-Inger Ben Gurion University, Beer Sheva, 8410501, Israel                        Bracha Shapira                        Lior Rokach
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing tasks  Chang et al. (2024); Matarazzo and Torlone (2025). Yet, when confronted with complex multitask benchmarks such as the Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020), a single LLM often struggles to excel uniformly across all subjects. The MMLU benchmark encompasses diverse subjects, each varying in difficulty and domain breadth. These variations pose significant challenges, revealing performance gaps that no individual model can easily bridge.
Ensembling multiple LLMs provides a promising avenue to overcome these limitations Tekin et al. (2024); Jiang et al. (2023b); Mavromatis et al. (2024); Xu et al. (2024b). By leveraging the complementary strengths of different models, an ensemble can achieve greater accuracy, robustness, and adaptability than any single model alone Lu et al. (2024). However, maximizing the potential of an ensemble requires careful selection and integration of its models. Key challenges include identifying models that offer complementary strengths, preserving a diverse set of solution strategies, adapting to subject-specific difficulties, and efficiently aggregating predictions.
Table 1: Comparison of related methods. “Zero/Few-Shot Training” indicates methods that do not require additional training or fine-tuning. DFPE requires no additional training, uses few-shot validation for subject adaptivity, and ensures diversity via clustering and quantile-based filtering.
Method Zero/Few-Shot Training Subject Adaptivity Diversity Optimization
LLM-TOPLA ×
SelectLLM × ×
PackLLM ×
SweetSpan ×
DeePEn ×
EVA ×
Boosted Prompts ×
CAPE ×
ZOOTER ×
LoRA Ensembles × ×
DFPE (Ours)
In this paper, we present the Diverse Fingerprint Ensemble (DFPE) method that aims to optimize the ensemble for a particular subject, thereby improving the performance of LLMs on that subject. Our approach optimizes the ensemble of models by combining effective model selection, subject-level adaptivity, and adaptive weighting. The key contributions of our method include:
  • Diversity Preservation through Response Pattern Clustering: To ensure the ensemble incorporates diverse problem-solving strategies, we capture each model’s response patterns through validation-set "fingerprints" and cluster them using DBSCAN. This systematically maintains complementary approaches while filtering redundancy, preventing the ensemble from converging to a single solution path and enhancing its ability to handle varying question types.
  • Subject-Specialized Expertise Allocation: To balance domain-specific capabilities with strategic diversity, we implement a two-stage process: first clustering models with similar response patterns, then selecting the highest-performing model from each cluster based on subject-specific validation accuracy. These selected models receive weights scaled by their subject performance, creating an ensemble that combines diverse approaches while emphasizing proven expertise.
  • Adaptive Performance Thresholding: To maintain quality while preserving valuable minority perspectives, we employ subject-level quantile-based filtering that automatically adjusts to varying difficulty levels across disciplines. This approach ensures high standards while accommodating the inherent differences between subjects, retaining models that might excel on specific question types.
We evaluate our approach on the MMLU benchmark using a diverse pool of up to 9 billion parameters LLMs. Our ensemble achieves a final accuracy of 73.5%, outperforming the best single model by 3%, and improves the discipline-level accuracy by 5%. These results demonstrate that our approach, which prioritizes selecting the most diverse and accurate LLMs tailored to each subject, leads to significant improvements in performance on complex, multi-faceted language understanding tasks.
Figure 1: A pool of LLMs $\mathcal{M}$ is evaluated per subject $S_{k}\in\mathcal{S}$ using a small validation set $\mathcal{Q}_{k}$. Each model’s predictions and accuracy $\alpha_{i,k}$ are used to generate “fingerprints,” which are subsequently clustered to maintain diversity. Models failing to meet a subject-specific quantile threshold are removed, and the most accurate model is chosen from each cluster to form the representative set $\mathcal{M}_{k}^{*}$. An exponential weighting scheme is then applied to these representatives before their final, weighted votes are aggregated to produce the answer
2 Related Work
Ensembling techniques have been widely explored as a means to boost the performance of LLMs on complex multitask benchmarks like MMLU. These methods typically aim to balance improved accuracy with practical constraints such as computational overhead, training requirements, and the diversity of model contributions. Some strategies focus on maximizing diversity across model outputs to exploit complementary strengths, while others emphasize efficient routing or fine-tuning to adapt ensemble components.
A prominent direction involves assembling multiple models to capitalize on their unique abilities, as seen in LLM-TOPLA Tekin et al. (2024), PackLLM Mavromatis et al. (2024), and SweetSpan Xu et al. (2024b), each of which coordinates output information at various granularities (e.g., token- or span-level). While these techniques often produce strong results, they may also introduce extra training or inference overhead. Similar concepts appear in ZOOTER Lu et al. (2023), which dynamically routes queries to suitable models, and LoRA Ensembles Wang et al. (2023), which uses parameter-efficient fine-tuning to merge model capabilities. However, approaches that rely heavily on additional training or fine-tuning can be less practical when resources are limited or when rapid experimentation is required.
Other methods seek to remain fully or mostly training-free. Boosted Prompts Pitis et al. (2023) and CAPE Jiang et al. (2023c), for instance, forgo further training in favor of leveraging pretrained models and calibrating their outputs. This philosophy aligns with EVA Xu et al. (2024a) and DeePEn Huang et al. (2024) in that they preserve model diversity or transform model outputs at an inference or validation stage without extensive fine-tuning. However, many of these solutions do not adapt well to subject-wise performance variations, which becomes crucial for wide-ranging benchmarks like MMLU.
SelectLLM Maurya et al. (2024) provides an example of per-query adaptivity, where a routing scheme decides which model should handle a particular input. Yet, these methods may not systematically ensure the retention of diverse solution paths or fine-grained subject-level adaptivity.
Table 1 compares these ensemble approaches with respect to (a) zero/few-shot training needs, (b) subject adaptivity, and (c) explicit diversity optimization. As shown, many existing solutions either require training or lack robust subject-wise specialization. Our method, DFPE, is distinguished by its reliance on few-shot validation rather than intensive training, its explicit per-subject clustering and filtering to maintain competence and diversity, and an adaptive weighting mechanism that emphasizes strong models without discarding minor yet potentially useful perspectives. This combination ensures that DFPE remains both robust and subject adaptive, providing a substantial improvement on language understanding tasks without demanding additional fine-tuning.
Overall, DFPE stands out by forgoing further training or complex fine-tuning, relying instead on validation-driven selection and weighting to construct a subject-adaptive, diversity-optimized ensemble that excels on challenging multitask benchmarks like MMLU.
3 Method
In this section, we detail our ensemble method designed to optimize LLMs ensemble for a specific subject tasks. Our approach integrates clustering-based selection, quantile-based filtering, and adaptive weighting to ensure that only sufficiently capable and diverse models contribute to the final ensemble prediction. By leveraging per-subject validation performance and subject-specific fingerprints, we construct a more balanced and effective ensemble than naive averaging or manual heuristics.
3.1 Overview
As illustrated in Figure 1, our method involves four main steps:
  1. 1.
    Model Fingerprinting and Clustering: Represent each model’s responses on a subject as a fingerprint vector. Cluster these fingerprints using DBSCAN with cosine similarity, grouping models exhibiting similar response patterns.
  2. 2.
    Quantile-based Model Selection: Identify a quantile-based accuracy threshold per subject and retain only models meeting or exceeding this benchmark. From each cluster, select the model with the highest accuracy, to ensure both competence and diversity.
  3. 3.
    Adaptive Weighting Scheme: Assign weights to the selected models using exponential scaling of their per-subject validation accuracy. Normalize these weights to form a proper distribution.
  4. 4.
    Ensemble Prediction: Aggregate the weighted predictions of the selected models via a weighted voting mechanism, producing the final answer.
This pipeline ensures that the ensemble benefits from both model diversity and subject-specific adaptivity, with performance-driven weighting giving more influence to stronger models.
3.2 Formal Description
Notation
  • Let =$\mathcal{M}=\{M_{1},M_{2},\dots,M_{N}\}$ be the set of available LLMs.
  • Let $\mathcal{S}=\{S_{1},S_{2},\dots,S_{K}\}$ represent the subjects in MMLU.
  • For each subject $S_{k}$, let $\mathcal{Q}_{k}=\{q_{k,1},q_{k,2},\dots,q_{k,L_{k}}\}$ denote its questions.
  • Each question $q_{k,l}$ has choices $\mathcal{C}_{k,l}=\{c_{k,l,1},\dots,c_{k,l,C_{k,l}}\}$.
  • Let $\hat{c}_{i,k,l}$ be the prediction of model $M_{i}$ for question $q_{k,l}$.
  • Let $\alpha_{i,k}$ denote the validation accuracy of model $M_{i}$ on subject $S_{k}$.
Model Fingerprinting and Clustering
For each model $M_{i}$ and subject $S_{k}$, we produce a fingerprint vector $\mathbf{f}_{i,k}$ summarizing the model’s response behavior on validation data. For example, we can embed each output (e.g., using a pre-trained sentence embedding model) and average the embeddings to create the fingerprint. Once we obtain $\{\mathbf{f}_{i,k}\}_{i=1}^{N}$, we apply DBSCAN clustering with cosine similarity to group models into clusters based on their response patterns.
Quantile-based Model Selection
Define a quantile parameter $q$ (e.g., $q=0.1$) to set a subject-specific accuracy threshold. Let $\{\alpha_{i,k}\}_{i=1}^{N}$ be the accuracies of all models on $S_{k}$. Compute the $q$-quantile accuracy threshold $\alpha_{q,k}$ for subject $S_{k}$. Retain models satisfying $\alpha_{i,k}\geq\alpha_{q,k}$, ensuring a baseline level of competence per subject.
From the filtered models, select the representative model from each cluster:
$M^{*}_{k,j}=\arg\max_{M_{i}\in C_{j}}\alpha_{i,k}.$
The set of all such representatives for subject $S_{k}$ is $\mathcal{M}^{*}_{k}$.
Adaptive Weighting Scheme
For each selected model $M_{i}\in\mathcal{M}^{*}_{k}$, we assign a weight:
$w_{i,k}=\exp(\gamma\cdot\alpha_{i,k}),$
where $\gamma$ is a scaling factor that highlights performance differences. Normalize these weights so that they sum to 1:
$\tilde{w}_{i,k}=\frac{w_{i,k}}{\sum_{M_{j}\in\mathcal{M}^{*}_{k}}w_{j,k}}.$
Ensemble Prediction
For a test question $q_{k,l}$ in subject $S_{k}$, each selected model $M_{i}\in\mathcal{M}^{*}_{k}$ provides a predicted choice $\hat{c}_{i,k,l}$. We aggregate their votes using the assigned weights:
$V_{k,l}(c)=\sum_{M_{i}\in\mathcal{M}^{*}_{k}}\tilde{w}_{i,k}\cdot\mathbb{I}[\hat{c}_{i,k,l}=c].$
The final ensemble prediction is:
$\hat{c}_{k,l}=\arg\max_{c\in\mathcal{C}_{k,l}}V_{k,l}(c).$
Algorithm 1 DFPE (Diverse Fingerprint Ensemble)
for each subject $S_{k}$ do
     1. For models $M_{i}$, compute accuracy $\alpha_{i,k}$ on validation $\mathcal{Q}_{k}$, filter out those below $q$-quantile.
     2. Create fingerprints $\mathbf{f}_{i,k}$ (e.g., embeddings), cluster via DBSCAN.
     3. From each cluster, pick the model $M_{i}$ with highest $\alpha_{i,k}$; call this set $\mathcal{M}_{k}^{*}$.
     4. Assign weights $w_{i,k}=\exp(\gamma\alpha_{i,k})$ to models in $\mathcal{M}_{k}^{*}$ and normalize.
end for
for each test question $q_{k,l}$ in $S_{k}$ do
     Aggregate predictions with weighted votes:
$\hat{c}_{k,l}=\arg\max_{c}\sum_{M_{i}\in\mathcal{M}_{k}^{*}}\bigl{(}w_{i,k}\cdot\mathbb{I}[\hat{c}_{i,k,l}=c]\bigr{)}.$
end for
3.3 Rationale
Our method strikes a balance between diversity, competence, and adaptability:
  • Diversity Preservation: Clustering models ensures variety in solution strategies, reducing redundancy.
  • Quantile-based Competence: The quantile threshold removes weak models, guaranteeing a baseline performance level.
  • Adaptive Weighting: Higher accuracy models naturally exert greater influence, but all retained models still contribute to the final decision.
  • Practical Simplicity: Operating at the output level avoids complex token- or span-level computations, making the method more scalable.
By combining these elements, our approach effectively exploits the complementary strengths of multiple LLMs, yielding improved performance on complex multitask language understanding.
4 Experimental Setup
In this section, we detail the dataset, base models, and implementation specifics used to evaluate our improved ensemble method. We rely on the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2020), incorporate a set of Large Language Models (LLMs) with varying capabilities, and use few-shot validation to guide quantile-based selection and adaptive weighting. We present the final results alongside attached figures illustrating discipline-level performance, sensitivity analyses, and accuracy comparisons.
4.1 Dataset
We evaluate our method on the MMLU benchmark Hendrycks et al. (2020), which consists of 57 subjects spanning a diverse range of disciplines, including STEM fields, humanities, and social sciences. Each subject comprises multiple-choice questions that vary in difficulty. The given benchmark data were partitioned into a few-shot validation set for model selection and fingerprint generation, and a separate test set for the final evaluation.
Table 2: Overall accuracies of the base LLMs on the MMLU benchmark. The diverse range of performance provides a strong foundation for our quantile-based ensemble approach.
Model Params Accuracy
GLM-4 9B 0.7076
Qwen2.5 7B 0.6476
Qwen2.5 3B 0.6605
Mistral v0.3 7B 0.6316
Phi-3.5-mini 4B 0.7046
Llama-3.1 8B 0.6508
Gemma-2 9B 0.6804
Apollo2 7B 0.7034
Starling-LM-alpha 7B 0.6149
Yi-1.5 6B 0.6239
4.2 Base Models
Our ensemble is constructed from a pool of LLMs differing in architecture, size, and training methodology, including GLM-4-9B-chat GLM et al. (2024), Qwen2.5-7B-Instruct Yang et al. (2024), Qwen2.5-3B-Instruct Yang et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023a), Phi-3.5-mini-instruct Abdin et al. (2024), Llama-3.1-8B-Instruct Dubey et al. (2024), Gemma-2-9B Team (2024), Apollo2-7B Zheng et al. (2024), Starling-LM-7B-alpha Zhu et al. (2023), and Yi-1.5-6B Young et al. (2024). These models are detailed in Table 2. The best single model in our experiments achieves 70.76% overall-accuracy on MMLU, with a discipline-accuracy of 69%. Other models range from approximately 61.5% to 70.5% overall-accuracy, providing a rich diversity of performance levels. This variability creates an opportunity to exploit complementary strengths: by carefully selecting and weighting models via quantile-based filtering and clustering, we aim to surpass the capabilities of any single model.
4.3 Implementation Details
Few-Shot Validation for Selection and Weighting:
For each subject, given validation questions (approximately 10% of the test set) to measure model accuracy. These validation accuracies guide the quantile threshold, clustering-based selection, and the adaptive weighting scheme. By utilizing this few-shots validation approach rather than additional training or fine-tuning, our method maintains computational efficiency while being compatible with any pre-trained LLM, whether accessed through local deployment or API endpoints.
Clustering and Quantile Thresholding:
For each subject, we embed each model’s validation responses to form a fingerprint vector. We then cluster these fingerprints using DBSCAN with cosine similarity. The DBSCAN epsilon parameter is chosen based on empirical tuning to ensure stable, meaningful clusters. We select a quantile parameter $q$ = 0.05 based on initial validation performance. Models meeting or exceeding the $q$-quantile accuracy threshold are retained. This filtering guarantees that only models demonstrating a baseline level of competence for that subject remain candidates for the ensemble.
Adaptive Weighting and Hyperparameters:
We apply exponential scaling to each retained model’s per-subject validation accuracy. Parameters such as the AccuracyFactor and quantile thresholds are determined via validation-based exploration. For example, setting the AccuracyFactor ($\gamma$) to 5.0 and choosing $q=0.05$ consistently yields strong performance. After weighting models according to their accuracies, we aggregate their predictions with a weighted voting mechanism, producing final ensemble predictions without extra fine-tuning or complex training procedures.
Reproducible Code
The source code and raw experiment results are available at https://github.com/nivgold/DFPE.
4.4 Evaluation Metrics
We report both the Overall Accuracy and Discipline-Accuracy which accounts for discipline-level balance, ensuring improvements are not limited to a few disciplines. We compare four configurations:
  • Best Single Model (BSM): The single model achieving the highest test accuracy is which accounts for subject-level balance, ensuring improvements are not limited to a few subjects.
  • Best Single Model on Validation (BSMoV): The model that performs best on the validation set.
  • Majority Voting Ensemble (MVoting): Equal weights ensemble method.
  • Our Ensemble (DFPE): The proposed quantile-driven cluster-based selection method with adaptive weighting is named DFPE.
These comparisons demonstrate how DFPE can outperform all the compared methods.
5 Results
We evaluate the performance of DFPE through extensive experiments, analyzing overall accuracy improvements, discipline-specific gains, and the impact of various hyperparameters. To enhance clarity, we aggregated the results at the subject level into discipline-level outcomes. Our findings demonstrate significant and consistent improvements over baseline approaches across a wide range of disciplines.
5.1 Overall Performance
DFPE consistently outperforms all comparison methods on the MMLU benchmark. As shown in Table 3, it achieves higher overall accuracy and discipline-specific accuracy than the best single models (BSM, BSMoV) as well as majority voting (MVoting). These results underscore the effectiveness of DFPE’s design, which leverages quantile-based filtering to identify top-performing predictions, clustering to preserve diverse model insights, and adaptive weighting to integrate complementary strengths into a unified ensemble.
Table 3: Accuracy and Discipline-Accuracy Comparison. DFPE significantly outperforms the Best Single Model (BSM), the best Single Model on Validation (BSMoV), and the Majority Voting method (MVoting).
Model Accuracy Discipline-Accuracy
BSM 0.708 0.690
BSMoV 0.677 0.676
MVoting 0.724 0.727
DFPE (Optimal) 0.735 0.740
5.2 Discipline-level Analysis
Table 4 provides a detailed breakdown of our results across 20 distinct disciplines. Overall, DFPE attains the highest average accuracy (0.740) and achieves the top score in 9 out of 20 disciplines (Business, Chemistry, Communication, Computer Science, Engineering, Ethics, Logic, Physics, and Statistics). As visualized in Figure 2, DFPE consistently matches or outperforms competing methods across a diverse set of tasks. These results underscore the strength of DFPE’s diversity-preserving design, which allows it to integrate the complementary knowledge of multiple base models and deliver robust, broad-spectrum performance.
Table 4: Comparison of discipline-level Results
Discipline BSM BSMoV MVoting DFPE
Biology 0.747 0.734 0.777 0.773
Business 0.724 0.665 0.773 0.775
Chemistry 0.460 0.475 0.593 0.605
Com 0.583 0.609 0.709 0.718
Comp Sci 0.581 0.629 0.668 0.690
Economics 0.794 0.678 0.738 0.756
Engineering 0.500 0.683 0.703 0.717
Ethics 0.614 0.654 0.607 0.674
History 0.686 0.741 0.818 0.815
Humanities 0.844 0.781 0.851 0.847
Law 0.683 0.690 0.729 0.726
Logic 0.722 0.749 0.822 0.871
Math 0.489 0.462 0.468 0.481
Medicine 0.890 0.784 0.825 0.823
Misc 0.701 0.586 0.656 0.654
Other 0.864 0.766 0.804 0.810
Physics 0.585 0.590 0.615 0.650
Politics 0.878 0.822 0.856 0.868
Social Sci 0.880 0.810 0.847 0.848
Statistics 0.565 0.611 0.690 0.694
Average 0.690 0.676 0.727 0.740
Figure 2: Average Accuracy by Discipline. DFPE consistently outperforms the compared methods across a wide range of disciplines, highlighting broad-spectrum gains.
5.3 Sensitivity Analysis
We conduct comprehensive sensitivity analyses across three key hyperparameters, as visualized in Figure 3:
Quantile Threshold: Optimal performance at 0.05.
- Stable performance range: 0.05-0.2
- Performance degradation beyond 0.10
AccuracyFactor: Best results at 5.
- Effective range: 4.0-6.0
- Diminishing returns beyond 6.0
DBSCAN Epsilon: Robust clustering achieved with low epsilon.
- Stable performance: 0.0001-0.0005
- Higher epsilon values result in an insufficient number of clusters, and when the epsilon value exceeds 0.01, all models fall into the same cluster, yielding results equivalent to the best single model.
Figure 3: Sensitivity analysis. Left: Accuracy vs. Quantile Threshold; Middle: Accuracy vs. AccuracyFactor; Right: Accuracy vs. Epsilon - the Epsilon axis (log scale). Performance remains robust within moderate parameter ranges, easing the tuning process.
5.4 Model Participation and Efficiency
In our optimal accuracy configuration, DFPE typically retains most of the available models (9 out of 10) per subject. This suggests that in the highest-performing setting, the diversity and complementary strengths of nearly all models contribute to the ensemble’s success. Although embedding responses, clustering, and validation-based selection introduce some overhead, the overall method remains efficient and does not require additional training or fine-tuning. For scenarios where computational efficiency is prioritized over maximum accuracy, we provide alternative configurations in Appendix A that achieve a balance between performance and model count.
5.5 Discussion
Our extensive experimental results provide several key insights into the effectiveness and practical implications of DFPE. We organize our discussion around three main aspects: performance characteristics, trade-offs, and practical considerations.
Performance Characteristics
Our method’s strong results can be attributed to effectively leveraging a diverse set of models via clustering, quantile-based filtering, and adaptive weighting. By preserving model heterogeneity, DFPE capitalizes on complementary strengths, often retaining a majority of models for each subject while assigning higher influence to those that demonstrate subject-specific expertise. In addition, recent work has highlighted new ensembles like Mixtral-of-Experts (8x7B) from Mistral.ai [^1] https://mistral.ai/news/mixtral-of-experts/ that achieve strong performance across multitask benchmarks. While some novel LLMs can surpass DFPE in absolute accuracy, they generally demand far larger computational resources. Notably, DFPE still exceeds certain significantly larger models, including LLaMA-2 70B, reinforcing the effectiveness of its ensemble approach. Incorporating those newer, larger LLMs into the DFPE model pool represents a promising direction for future work, potentially unlocking even further gains in performance through expanded ensemble diversity.
Parameter Sensitivity and Trade-offs
Our sensitivity analyses, shown in Figure 3, reveal important trade-offs between accuracy, computational cost, and model participation:
Quantile Threshold: A low threshold (0.05) maximizes accuracy by including more models but increases computational cost. Higher thresholds (0.15-0.20) offer a more balanced trade-off, reducing the ensemble size while maintaining most performance benefits.
AccuracyFactor: Higher values (around 5.0) effectively emphasize stronger models’ contributions, but excessive values risk over-reliance on top performers, potentially reducing the benefits of ensemble diversity.
DBSCAN Epsilon: Smaller values produce more fine-grained clusters, preserving subtle model differences, while larger values lead to more aggressive grouping and smaller ensembles.
For practical applications, we identify three deployment configurations that balance performance and efficiency:
High-Accuracy Mode: Uses most available models, achieving maximum accuracy at higher computational cost. Recommended for applications where accuracy is paramount.
Balanced Mode: Employs higher quantile thresholds and larger clustering epsilon, reducing ensemble size to 5-7 models while maintaining performance within 0.5-1.0
Efficient Mode: Further reduces model count through more aggressive filtering, suitable for resource-constrained environments while still outperforming single models.
These configurations, detailed in Appendix A, provide practitioners with flexible options based on their specific requirements.
Limitations
While DFPE demonstrates significant improvements, several limitations warrant discussion:
Cold-start Performance: The method requires validation data for initial model selection and weighting, which may not always be available for new subjects or domains.
Computational Overhead: In its highest-accuracy configuration, the method requires running multiple models, which may be impractical for some real-time applications.
Model Diversity Requirement: The effectiveness of DFPE depends on having access to a diverse pool of models with complementary strengths.
6 Conclusion
In this paper, we introduced DFPE, an ensemble method that leverages diversity and adaptivity to improve LLMs performance on multitask language understanding tasks. DFPE clusters models via “fingerprint” patterns, filters underperformers with a quantile-based threshold, and applies exponential weighting that emphasizes top-performing models while preserving valuable secondary perspectives. On the MMLU benchmark, DFPE achieves approximately 3% higher overall accuracy and a 5% boost in discipline-level accuracy over the best single model. These results underscore the importance of diverse solution strategies, selective filtering, and dynamic weighting. While DFPE currently focuses on multiple-choice tasks, applying its fingerprinting strategy to open-ended settings remains a promising direction. Future work also includes refining question-level adaptivity, exploring more scalable clustering, and addressing the unique challenges posed by large LLM pools. By balancing diversity, adaptivity, and efficiency, DFPE offers a practical and robust framework for high-performing ensembles across varied tasks.
Appendix A Appendix A
While our main results focus on the optimal accuracy configuration, practitioners often need to balance performance gains against computational costs. Here we present a detailed analysis of an alternative configuration that achieves near-optimal performance while significantly reducing computational overhead.
A.1 Balanced Configuration Parameters
We identified a balanced configuration with the following parameters:
  • Quantile threshold: 0.5 (vs. 0.05 in optimal setting)
  • AccuracyFactor: 7 (vs. 5 in optimal setting)
  • DBSCAN Epsilon: 0.001 (vs. 0.0001 in optimal setting)
This configuration achieves an average accuracy of 72.5% (1% below optimal) while reducing the mean number of models per subject to 6 (compared to 9 in the optimal setting).
A.2 Model Selection Analysis
Figure 4 shows the distribution of selected models across subjects. Several key patterns emerge:
  • Variation in model count: The number of selected models varies from 1 to 10 across subjects
  • Subject-specific adaptation: Different subjects benefit from different ensemble sizes
  • Consistent core: Most subjects maintain 6-8 models, suggesting a natural balance point
Figure 4: Distribution of selected models per subject. The variation in bar heights demonstrates how DFPE adapts its ensemble size to subject-specific requirements while maintaining efficiency.
A.3 Model Co-occurrence Analysis
To understand model relationships, we analyzed their co-occurrence patterns within clusters (Figure 5). Our analysis reveals several interesting patterns:
  • Strong Partnerships: - Qwen family models (Qwen2.5-3B and Qwen2.5-7B-Instruct) show highest co-occurrence (52 instances) - Strong affinity between Qwen models and Llama-3.1-8B (51-52 co-occurrences) - Phi-3.5-mini frequently pairs with Qwen models (48-49 instances)
  • Complementary Groups: - Models cluster into "specialists" and "generalists" - Lower co-occurrence patterns indicate complementary strengths
  • Model Independence: - Some models show consistent independence - Suggests unique capabilities or specialization
Figure 5: Heatmap of model co-occurrences within clusters. Cell values indicate frequency of model pairs being selected together. The diagonal is zero by definition. Higher values (darker colors) suggest stronger complementarity between models.
A.4 Practical Implications
These findings have several important implications for practitioners:
Resource Optimization: The balanced configuration offers a practical trade-off between performance and computational cost
Model Selection: Strong co-occurrence patterns can guide initial model selection when building new ensembles
Deployment Strategy: Subject-specific ensemble sizes suggest opportunities for dynamic resource allocation